Here is my question.
You have a rectangle with 78 square centimeters.What are the dimensions of the rectangle with the largest perimeter?
Im not so sure about how to do this. Please help!

Answer by stanbon(75887)

You have a rectangle with 78 square centimeters.What are the dimensions of the rectangle with the largest perimeter?
------
Equation:
LW = 78 sq. cm
Perimeter = 2(L + W)
---
P = 2(L + (78/L))
---
P = 2(L^2+78)/L
---
P = (2L^2+156)/L
---
Take the derivative:
P' = [L(4L)-(2L^2+156)]/L^2
------
P' = [2L^2-156]/L^2
-----
Solve: P' = 0
---
2L^2-156 = 0
L^2 = 78
Length = sqrt(78)
-----
Width = sqrt(78)
=======================
Cheers,
Stan H.
===============